# Quantum field theory treatment of oscillations of Dirac neutrinos in external fields

### v1

@@ Appendix A Propagators of Dirac neutrinos in matter
where $m_{a}$ is the mass of $\psi_{a}$, $g_{a}$ is the diagonal effective potential of the matter interaction, $\bm{\alpha}=\gamma^{0}\bm{\gamma}$ and $\beta=\gamma^{0}$ are the Dirac matrices, and $\mathbf{p}=-\mathrm{i}\nabla$ is the momentum operator. Equation (A.1) can be obtained from Eqs. 2.3 and (4.1) if we omit there the nondiagonal potential, $g=0$.
@@
$\left\{a_{a\sigma}(\mathbf{p}),a_{a\sigma^{\prime}}^{\dagger}(\mathbf{q})\right\}=\left\{b_{a\sigma}(\mathbf{p}),b_{a\sigma^{\prime}}^{\dagger}(\mathbf{q})\right\}=\delta_{\sigma\sigma^{\prime}}\delta(\mathbf{p}-\mathbf{q}),$ (A.11)
with the rest of anticommutators being equal to zero, one gets that $\left\{\psi(\mathbf{x},t),\psi^{\dagger}(\mathbf{y},t)\right\}=\delta(\mathbf{x}-\mathbf{y})$. Thus, $\psi_{a}$ in Eq. (A.2) is the properly quantized solution of Eq. (A.1).
@@
$\times\bigg[\left(m_{a}+(\sigma p-g_{a}/2)\gamma^{5}\gamma^{0}\right)\left(\frac{1}{p_{0}-E_{a\sigma}-g_{a}/2+\mathrm{i}0}-\frac{1}{p_{0}+E_{a\sigma}-g_{a}/2-\mathrm{i}0}\right)$
@@
$\gamma^{0}\otimes\bm{\sigma}\equiv\left(\begin{array}[]{cc}0&-\bm{\sigma}\\ -\bm{\sigma}&0\end{array}\right),$ (A.14)
@@
$S_{a}(p_{0},\mathbf{p})=-\frac{1}{4(p_{0}-E_{a-}-g_{a}/2+\mathrm{i}0)}(1-\alpha_{a}\gamma^{5})\left(\begin{array}[]{cc}0&1-\xi\left(\bm{\sigma}\hat{p}\right)\\ 1-\xi\left(\bm{\sigma}\hat{p}\right)&0\end{array}\right),$ (A.16)
@@ Appendix B Computation of integrals
The integration over $z$ is characterized by the poles, which are the roots of the equation $E-\mathcal{E}_{+}+\mathrm{i}0=0$. One can show that these roots are in the upper half-plane. Nevertheless, it is not an easy task to find all the roots analytically. We find an approximate solution of this equation for ultrarelativistic neutrinos. The roots are $z_{+}=\sqrt{\rho_{0}^{2}-\rho^{2}}$ if $\rho<\rho_{0}$, and $z_{+}=\mathrm{i}\sqrt{\rho^{2}-\rho_{0}^{2}}$ if $\rho>\rho_{0}$, where $\rho_{0}=\bar{E}_{m}+\mathfrak{E}_{m}$, with $\bar{E}_{m}$ and $\mathfrak{E}_{m}$ given in Eq. (4.28).
@@
$=-\frac{2\mathrm{i}e^{\mathrm{i}\rho_{0}L}}{L}\left(1+\frac{\mathrm{i}}{2\rho_{0}L}\right)\approx-\frac{2\mathrm{i}e^{\mathrm{i}\rho_{0}L}}{L},$ (B.3)
where we assume that the propagation length is long enough, $L\gg\bar{E}_{m}^{-1}$.
Equation (B.3) allows one to finalize the computation of $I_{+}$ in Eq. (B.2). The expression for $I_{-}$ can be found analogously. We provide just the final results for both integrals,
$I_{\pm}=-\frac{\bar{E}_{m}e^{\mathrm{i}(\bar{E}_{m}\pm\mathfrak{E}_{m})L}}{4\pi L}\frac{\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)\sin 2\theta+g\cos 2\theta}{\sqrt{\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)^{2}+g^{2}}}.$ (B.4)
@@
$z_{+}=\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}_{\mathrm{B}}}\cos^{2}\phi}\times\begin{cases}\sqrt{\rho_{0}^{2}-\rho^{2}},&\text{if}\ \rho<\rho_{0},\\ \mathrm{i}\sqrt{\rho^{2}-\rho_{0}^{2}},&\text{if}\ \rho>\rho_{0},\end{cases}$ (B.8)
@@
Calculating the integral over $\rho$ in Eq. (B.7) analogously to Eq. (B.3) by splitting the integration segment as $0,\infty)=[0,\rho_{0}]\cup[\rho_{0},\infty)$, one gets that
@@
Then, we assume that $L\gg E^{-1}$ and calculate the remaining integral over $\phi$ in Eq. (B) in the approximation $(\mu B)^{2}\ll E\mathfrak{E}_{\mathrm{B}}$. We present the final result for both $I_{+}$ and $I_{-}$, which is calculated analogously,
$I_{\pm}=-\frac{Ee^{\mathrm{i}(E\pm\mathfrak{E}_{\mathrm{B}})L}}{4\pi L}\frac{\sin 2\theta\frac{\Delta m^{2}}{4E}}{\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}}\left[1-\frac{(\mu B)^{2}}{2E\mathfrak{E}_{\mathrm{B}}}\right].$ (B.11)
We use Eq. (B.11) in Eq. (5.35).
@@ Appendix C Propagators of Dirac neutrinos in vacuum
$S_{a}(x)=$ $\frac{1}{2}\int\frac{\mathrm{d}^{4}p}{(2\pi)^{4}}e^{-\mathrm{i}px}\bigg[\frac{1}{p_{0}-E_{a}+\mathrm{i}0}\left(\gamma^{0}-\frac{p}{E_{a}}\bm{\gamma}\hat{p}+\frac{m_{a}}{E_{a}}\right)$
$+\frac{1}{p_{0}+E_{a}-\mathrm{i}0}\left(\gamma^{0}+\frac{p}{E_{a}}\bm{\gamma}\hat{p}-\frac{m_{a}}{E_{a}}\right)\bigg],$ (C.1)
@@
$S_{a}(p_{0},\mathbf{p})\to\frac{\gamma^{0}-\varsigma_{a}\bm{\gamma}\hat{p}}{2(p_{0}-E_{a}+\mathrm{i}0)},$ (C.2)

### v2

@@ Appendix A Propagators of Dirac neutrinos in matter
where $m_{a}$ is the mass of $\psi_{a}$, $g_{a}$ is the diagonal effective potential of the matter interaction, $\bm{\alpha}=\gamma^{0}\bm{\gamma}$ and $\beta=\gamma^{0}$ are the Dirac matrices, and $\mathbf{p}=-\mathrm{i}\nabla$ is the momentum operator. Equation (A.1) can be obtained from Eqs. (2.3) and (4.1) if we omit there the nondiagonal potential, $g=0$.
@@
$\left\{a_{a\sigma}(\mathbf{p}),a_{a\sigma^{\prime}}^{\dagger}(\mathbf{q})\right\}_{+}=\left\{b_{a\sigma}(\mathbf{p}),b_{a\sigma^{\prime}}^{\dagger}(\mathbf{q})\right\}_{+}=\delta_{\sigma\sigma^{\prime}}\delta(\mathbf{p}-\mathbf{q}),$ (A.11)
with the rest of anticommutators being equal to zero, one gets that $\left\{\psi(\mathbf{x},t),\psi^{\dagger}(\mathbf{y},t)\right\}_{+}=\delta(\mathbf{x}-\mathbf{y})$. Thus, $\psi_{a}$ in Eq. (A.2) is the properly quantized solution of Eq. (A.1).
@@
$\times\bigg[\left(m_{a}-\mathrm{i}(\sigma p-g_{a}/2)\sigma_{2}\right)\left(\frac{1}{p_{0}-E_{a\sigma}-g_{a}/2+\mathrm{i}0}-\frac{1}{p_{0}+E_{a\sigma}-g_{a}/2-\mathrm{i}0}\right)$
@@
$\sigma_{1}\otimes\bm{\sigma}\equiv\left(\begin{array}[]{cc}0&\bm{\sigma}\\ \bm{\sigma}&0\end{array}\right),$ (A.14)
@@
$S_{a}(p_{0},\mathbf{p})=-\frac{1}{4(p_{0}-E_{a-}-g_{a}/2+\mathrm{i}0)}(1-\alpha_{a}\gamma^{5})\left(\begin{array}[]{cc}0&1-\xi\left(\bm{\sigma}\hat{\mathbf{p}}\right)\\ 1-\xi\left(\bm{\sigma}\hat{\mathbf{p}}\right)&0\end{array}\right),$ (A.16)
@@ Appendix B Computation of integrals
The integration over $z$ is characterized by the poles, which are the roots of the equation $E-\mathcal{E}_{+}+\mathrm{i}0=0$. One can show that these roots are in the upper half-plane. Nevertheless, it is not an easy task to find all the roots analytically. We find an approximate solution of this equation for ultrarelativistic neutrinos. The roots are $z_{+}=\sqrt{\rho_{0}^{2}-\rho^{2}}+\mathrm{i}0$ if $\rho<\rho_{0}$, and $z_{+}=\mathrm{i}\sqrt{\rho^{2}-\rho_{0}^{2}}$ if $\rho>\rho_{0}$, where $\rho_{0}=\bar{E}_{m}+\mathfrak{E}_{m}$, with $\bar{E}_{m}$ and $\mathfrak{E}_{m}$ given in Eq. (4.28).
@@
$=-\frac{2\mathrm{i}e^{\mathrm{i}\rho_{0}L}}{L}\left(1+\frac{\mathrm{i}}{2\rho_{0}L}\right).$ (B.3)
To obtain Eq. (B) we use the values of the following definite integrals:
$\int_{0}^{\rho_{0}}\rho\mathrm{d}\rho\frac{e^{\mathrm{i}L\sqrt{\rho_{0}^{2}-\rho^{2}}}}{\sqrt{\rho_{0}^{2}-\rho^{2}}}$ $=-\frac{\mathrm{i}}{L}(e^{\mathrm{i}\rho_{0}L}-1),$
$\int_{0}^{\rho_{0}}\rho\mathrm{d}\rho e^{\mathrm{i}L\sqrt{\rho_{0}^{2}-\rho^{2}}}$ $=-\frac{\mathrm{i}}{L^{2}}e^{\mathrm{i}\rho_{0}L}(L\rho_{0}+\mathrm{i})-\frac{1}{L^{2}},$
$\int_{0}^{\rho_{0}}\rho^{3}\mathrm{d}\rho\frac{e^{\mathrm{i}L\sqrt{\rho_{0}^{2}-\rho^{2}}}}{\sqrt{\rho_{0}^{2}-\rho^{2}}}$ $=-\frac{2}{L^{3}}e^{\mathrm{i}\rho_{0}L}(L\rho_{0}+\mathrm{i})+\frac{\mathrm{i}}{L^{3}}(L^{2}\rho_{0}^{2}+2),$
$\int_{\rho_{0}}^{\infty}\rho\mathrm{d}\rho\frac{e^{-L\sqrt{\rho^{2}-\rho_{0}^{2}}}}{\sqrt{\rho^{2}-\rho_{0}^{2}}}$ $=\frac{1}{L},$
$\int_{\rho_{0}}^{\infty}\rho\mathrm{d}\rho e^{-L\sqrt{\rho^{2}-\rho_{0}^{2}}}$ $=\frac{1}{L^{2}},$
$\int_{\rho_{0}}^{\infty}\rho^{3}\mathrm{d}\rho\frac{e^{-L\sqrt{\rho^{2}-\rho_{0}^{2}}}}{\sqrt{\rho^{2}-\rho_{0}^{2}}}$ $=\frac{2}{L^{3}}+\frac{\rho_{0}^{2}}{L}.$ (B.4)
As a rule, one assumes that the propagation length is long enough, $L\gg\bar{E}_{m}^{-1}$. However, we keep the terms $\propto L^{-2}$ to study the behavior of the probabilities at relatively short distances.
Equation (B) allows one to finalize the computation of $I_{+}$ in Eq. (B.2). The expression for $I_{-}$ can be found analogously. We provide just the final results for both integrals,
$I_{\pm}=-\frac{\bar{E}_{m}e^{\mathrm{i}(\bar{E}_{m}\pm\mathfrak{E}_{m})L}}{4\pi L}\frac{\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)\sin 2\theta+g\cos 2\theta}{\sqrt{\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)^{2}+g^{2}}}\left(1+\frac{\mathrm{i}}{2(\bar{E}_{m}\pm\mathfrak{E}_{m})L}\right).$ (B.5)
@@
$z_{+}=\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}_{\mathrm{B}}}\cos^{2}\phi}\times\begin{cases}\sqrt{\rho_{0}^{2}-\rho^{2}}+\mathrm{i}0,&\text{if}\ \rho<\rho_{0},\\ \mathrm{i}\sqrt{\rho^{2}-\rho_{0}^{2}},&\text{if}\ \rho>\rho_{0},\end{cases}$ (B.9)
@@
The calculation of the integral over $\rho$ in Eq. (B.8) is analogous to Eq. (B). We split the integration segment as $0,\infty)=[0,\rho_{0}]\cup[\rho_{0},\infty)$ and use Eq. (B). Eventually, one gets that
@@
Then, we calculate the remaining integral over $\phi$ in Eq. (B) in the approximation $(\mu B)^{2}\ll E\mathfrak{E}_{\mathrm{B}}$. We present the final result for both $I_{+}$ and $I_{-}$, which is calculated in a similar manner,
$I_{\pm}=-\frac{Ee^{\mathrm{i}(E\pm\mathfrak{E}_{\mathrm{B}})L}}{4\pi L}\frac{\sin 2\theta\frac{\Delta m^{2}}{4E}}{\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}}\left(1-\frac{(\mu B)^{2}}{2E\mathfrak{E}_{\mathrm{B}}}\right)\left[1+\frac{\mathrm{i}}{2EL}\left(1-\frac{(\mu B)^{2}}{2\mathfrak{E}_{\mathrm{B}}^{2}}\right)\right].$ (B.12)
Analogously to Eq. (B.5), we keep only the lext-to-leading term $\propto L^{-2}$ in Eq. (B.12). We use Eq. (B.12) in Eq. (5.35).
@@ Appendix C Propagators of Dirac neutrinos in vacuum
$S_{a}(x)=$ $\frac{1}{2}\int\frac{\mathrm{d}^{4}p}{(2\pi)^{4}}e^{-\mathrm{i}px}\bigg[\frac{1}{p_{0}-E_{a}+\mathrm{i}0}\left(\gamma^{0}-\frac{p}{E_{a}}\bm{\gamma}\hat{\mathbf{p}}+\frac{m_{a}}{E_{a}}\right)$
$+\frac{1}{p_{0}+E_{a}-\mathrm{i}0}\left(\gamma^{0}+\frac{p}{E_{a}}\bm{\gamma}\hat{\mathbf{p}}-\frac{m_{a}}{E_{a}}\right)\bigg],$ (C.1)
@@
$S_{a}(p_{0},\mathbf{p})\to\frac{\gamma^{0}-\varsigma_{a}\bm{\gamma}\hat{\mathbf{p}}}{2(p_{0}-E_{a}+\mathrm{i}0)},$ (C.2)